A projectile is thrown at an angle theta Such that it is just able to cross a vertical wall at heighest point of journey the angle theta at which the projectile is thrown is given by?
1
Expert's answer
2019-05-13T10:32:52-0400

Maximum height is given by formula


"h=\frac {{v_0}^2 {\sinθ}^2}{2g} (1)"

Using (1) we get:


"θ=\sqrt{\arcsin{\frac {2gh}{{v_0}^2 }}}"